OpenStudy (anonymous):

Factor the polynomial: 32a^2b^2+16ab^2+2b^2 MEDAL WILL BE GIVEN TO THE BEST EXPLANATION!

OpenStudy (anonymous):

if u want to learn, at least try simplifying your equation a bit. by dividing everything by 2.

OpenStudy (anonymous):

or expression for that matter.

OpenStudy (anonymous):

wouldnt i divide by 4??

OpenStudy (mathmate):

Use grouping, but factor out the obvious common factors first: \( 32a^2b^2+16ab^2+2b^2 \\ =2b^2(16a^2+8a+1) \) Now factorize what's inside the parentheses, find m and n such that m+n=8, m*n=16*1=16 Proceed to grouping as in the other problem.

OpenStudy (anonymous):

last term is 2 so the gcf is 2

OpenStudy (anonymous):

@mathmate you can also factor out the a though?

OpenStudy (mathmate):

Sorry, but I don't see an \(a\) in the last term \(2b^2\), so \(a\) is not a common factor.

OpenStudy (anonymous):

@mathmate can you go over the "factorize inside the parentheses" part, you sorta lost me there

OpenStudy (mathmate):

Use grouping. Let ax^2+bx+c be the expression to be factorized. We look for integers m, n such that m+n=b, and m*n=a*c. For the given problem, a=16, b=8, c=1. So we look for m,n such that m+n=8, m*n=16*1=16 Can you find m, n satisfying these conditions?

OpenStudy (mathmate):

@ineedyouubiebs guess you're not online. I will explain the other steps, so you can continue when you're online. Here, we see that m=n=4 to give mn=16, m+n=8 So we group the expression as follows: 16x^2+4x + 4x + 1 =4x(4x+1) + (4x+1) take out the common factor (4x + 1) and you'll get =(4x+1)(4x+1) =(4x+1)^2 In this particular case, there is a shorter method: If both the x^2 term and the constant term are perfect squares, as in (4x)^2=16x^2, and 1^2 = 1 We double the product (1*4x) to get 8x and compare with the middle term. If they equal, then the expression is a perfect square, equal to (4x+1)^2.

OpenStudy (calculusxy):

http://library.thinkquest.org/20991/alg2/polyf.html